Prev Next

SW&IT Meeting Minutes 2021-09-27

Attendees: Brandon, David, Glenn, Matthieu, Paul, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20210927

LR20210920 - Bec, Matthieu 40.0
LR20210920 - Mechtley, Brandon 40.0
LR20210920 - Eychaner, Glenn 40.0
LR20210920 - Edmeades, David 40.0
LR20210920 - Grenz, Paul 12.0
LR20210920 - Hooper, Stephen 80.0
LR20210920 - Zhang, Yang 40.0
LR20210920 - Zhang, Xianyu 40.0

News from management

Previous Action Items

Owner Action Due Date Completion Date

Open topics

  • IDL for CV

Team reports

MB

  • Last week
    • OC, Management
    • dms m1surf fix unanimated updates
    • ATT PM
    • ao data/analysis discussion with BM
    • inotify discussion with YZ
    • IT8479 (radar offline), tcs-web wda.sh git update
    • IT expenditures mid Oct (after board meeting)
    • PG recruitment moving along 5-7 days
  • This week
    • git/github "policy" (= philosophy slides) for new staff
    • ao-analysis moving fwd
    • ao / att with BM
    • SWG reports, overdue
    • IT8470 pmc/thermal repeated telemetry

YZ

  • Previous:
    • Some work on new image notification program (some discussion with MB).
    • Work on MODS WCS keywords program.
    • Patched DMS to enable wcs client (wcs database collection now being populated).
    • TMS meeting, some discussion, and software work.
  • Next:
    • Continue work on new image notification program (and initiate some discussion with Al and Archive Group).
    • Continue to work on MODS WCS keywords program.
    • Some other misc MODS work.
    • Take a look at the DMS mongodb init program (and possibly extend it).
    • Continue some TMS work.
    • Work with Paul on Argos swingarm test/debug, if we get telescope time.

DE

  • VM clone work for BM, GE
  • Still nothing on the Windows server thing, followed up with Cindy this morning
  • Reprise going in to work on backup server
  • Various user account tasks
  • IssueTrak new feature: OAUTH integration. Discussed with IT, probably more trouble than it's worth.
  • Will prep vendor lists for Synology purchase so it's ready to go if/when we get the greenlight
  • Luci md125: for sure was a transient?

BM

  • Last week:
    • AO gain=0 test scripts, following testing/discussion.
    • Updated diagnostic dump parsing scripts to be used interactively.
    • Discussed AO data backup process with MB/XZ.
    • AO DD headers: delved into AO ICE implementation, WFS magnitude calculation.
  • This week:
    • I'm in Tucson now (new Zoom background).
    • Set up AO development VMs (for real this time).
      • Take care of ao-supervisor logging issues.
    • ao-analysis VM/data migration.
    • Continue following AO investigation.
    • Play more with AO ICE, getting DMS data (for ao-datasystems).

GE

  • Last week:
    • Went down a lot of rabbit holes, unfortunately.
    • Met with AO regarding LUCI image FITS headers and data dictionary.
    • Installed TCS and LUCI VMs copied by DE on my laptop.
      • Changed my home network from 10.0.0.x to 10.0.7.x so it doesn't overlap the telescope private network.
    • Started investigating and fixing the LUCI issues.
      • Checked everything out of GitHub with a proper fork-and-clone.
      • Changed from Eclipse to IntelliJ IDEA CE. The LUCI GitHub doesn't contain the project files for any IDE anyway.
  • This week:
    • Continue working on LUCI issues.

SH

  • Jim laptop, Mark software, Lenovo support
  • OVMS telemetry
  • UITS switches
  • VPN issues with Windows
  • UA risk management with Louise
  • LDAP
  • info page

next:

  • LDAP
  • IDL for Christian *
Topic revision: r11 - 27 Sep 2021, GlennEychaner
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback